Make up an equation of the form y = kx +b, the graph of which passes through the following points M (5, 5) and N (–10, –19)
Answer by rothauserc(4718)

We are given two points M(5, 5) and N(–10, –19)
:
k = slope of line = (-19-5)/(-10-5) = -24/-15 = 8/5
:
y = 8x/5 +b
:
use point M(5,5)
:
5 = 8(5)/5 +b
:
5 = 8 +b
:
b = -3
:
*************************
y = 8x/5 -3
*************************
: